OpenStudy (nottim):

Given the vector a=[-4.1,7], b=[2,0,-3], and c=[1,-1,5], simplify b*(a+c). Cartesian Vectors of a Vectors Unit.

OpenStudy (nottim):

My answer is [-6,0,-36]

OpenStudy (nottim):

Correct answer is -42.

OpenStudy (anonymous):

so, its b dot (a+c)?

OpenStudy (nottim):

Yes.

OpenStudy (anonymous):

a + c = [-3, 0, 12]

OpenStudy (anonymous):

[-3, 0, 12] dot [2, 0, -3] = -3(2) + 0(0) + -3(12) = -42

OpenStudy (nottim):

So I only add at the very end right?

OpenStudy (nottim):

Thanks.

OpenStudy (anonymous):

add a and b first, then take the dot product with c